Is there a solution to this?
Solve the equation and check your solution
0.40x + 0.20(x + 1) = 0.75

Answer by lwsshak3(11628)

Solve the equation and check your solution
0.40x + 0.20(x + 1) = 0.75
0.40x+.20x+0.20=0.75
.60x=.55
x=.55/.60=11/12
check:
0.40x + 0.20(x + 1)
=.40(11/12)+0.20(11/12+1)
=4.4/12+2.2/12+.2=6.6/12+0.2=0.55+0.2=0.75
check ok